OpenStudy (anonymous):

What is the slope of the line which passes through (1, –2) and (–1, 6)? 4 –4 2 –2

OpenStudy (anonymous):

let think we need to use the slope formula which is x2-x1/y2-y1

OpenStudy (anonymous):

then we use y=mx+b

OpenStudy (anonymous):

-1-1/6-(-2)=-2/8 simplify we get -1/4 cuz if we divide both the denominator and numer by 2 we get -1/4

OpenStudy (anonymous):

then we has the value for m=-1/4

OpenStudy (anonymous):

now we need to find the values of b

OpenStudy (anonymous):

u can pick either one (1, –2) or (–1, 6) i will use the first one which more easy to solve

OpenStudy (anonymous):

y=-2, x=1 so y=-1/4x+b -2=-1/4(1)+b cross mutiply -8/-1=b remember when we have neg and neg in the div or multi we get positve u final ans for finding he slope of the line which passes is y=-1/4x+8

OpenStudy (anonymous):

sorry my mistake is support to be y2-y1/x2-x1 opp side is -4 and the find the b is 2 which ur final answer is y=2x+2

OpenStudy (anonymous):

y=-4x+2

OpenStudy (anonymous):

i doing this too fast

OpenStudy (anonymous):

u ans should be 2